When 2x^3-ax^2+bx-c is divided by x+1 the remainder is -14
"determine algebraically the values for a, b an c"
I'm lost as to how to solve for a b and c

Answer by stanbon(48558)

When 2x^3-ax^2+bx-c is divided by x+1 the remainder is -14
"determine algebraically the values for a, b an c"
-----------------------
Use synthetic division as follows:
-1)....2.....-a.. ..b...........-c
........2......-2-a.....a+b+2.....|..-a-b-c-2
==============================
Remainder = -a-b-c-2 = -14
a + b+ c = 12
===============
Cheers,
Stan H.